# Evolutionary Generative Optimization: Towards Fully Data-Driven Evolutionary Optimization via Generative Learning

### v1

@@ Abstract
Recent advances in data-driven evolutionary algorithms (EAs) have demonstrated the potential of leveraging data to improve optimization accuracy and adaptability. Nevertheless, most existing approaches remain dependent on handcrafted heuristics, which limits their generality and automation. To address this challenge, we propose Evolutionary Generative Optimization (EvoGO), a fully data-driven framework empowered by generative learning. EvoGO streamlines the evolutionary optimization process into three stages: data preparation, model training, and population generation. The data preparation stage constructs a pairwise dataset to enrich training diversity without incurring additional evaluation costs. During model training, a tailored generative model learns to transform inferior solutions into superior ones. In the population generation stage, EvoGO replaces traditional reproduction operators with a scalable and parallelizable generative mechanism. Extensive experiments on numerical benchmarks, classical control problems, and high-dimensional robotic tasks demonstrate that EvoGO consistently converges within merely 10 generations and significantly outperforms a wide spectrum of optimization approaches, including traditional EAs, Bayesian optimization, and reinforcement learning based methods. Source code will be made publicly available.
Index Terms:
Generative learning, learning-based evolutionary algorithm, large-scale expensive optimization
@@ I Introduction
Many real-world optimization tasks involve objectives that are accessible only through simulations, legacy software, or physical experiments [1, 2]. Such complex problems are typically non-analytical and lack gradient information, presenting significant challenges for traditional optimization techniques. Evolutionary Algorithms (EAs) have emerged as a versatile and widely adopted framework for tackling such problems, with demonstrated success in domains such as hyperparameter tuning, biomedical engineering, and aerospace design [3]. Classical EAs maintain a population of candidate solutions and apply handcrafted variation operators (e.g., crossover and mutation) to iteratively explore the search space [4]. Although effective in many scenarios, these operators often struggle in high-dimensional or rugged landscapes, such as those found in complex robotic control tasks [5].
To enhance the performance of conventional EAs, researchers have introduced domain-specific heuristics into them. For instance, Evolutionary Strategies (ES)[6] and Estimation of Distribution Algorithms (EDAs)[7] employ statistical modeling to guide the search process. Despite their success in specialized contexts, these methods still rely on manually crafted rules and assumptions, thereby limiting their scalability, generality, and adaptability across diverse problem domains. With the increasing scale and complexity of modern engineering systems, along with the growing availability of data, conventional EAs face mounting challenges in navigating high-dimensional search spaces and addressing large-scale optimization tasks. This has led to a rising demand for more powerful EAs capable of handling increasingly sophisticated problems. To meet this demand, researchers have explored the integration of machine learning techniques into evolutionary computation, giving rise to a new class of learnable EAs [8]. These methods are generally categorized into two groups: non-data-driven and data-driven ones. The non-data-driven methods  [9, 10, 11, 12] typically enhance algorithmic components through predefined rules or online adaptation, but without systematically leveraging the data collected during the optimization process. In contrast, data-driven methods  [13, 14, 15, 16, 17] explicitly model and utilize the historical search data (e.g., evaluated solutions and their fitness values) to inform and guide future search decisions in a more principled and efficient manner.
A notable subclass within the data-driven category is generative modeling-based evolutionary algorithms (GMEAs) [18, 19, 20], which have recently attracted increasing attention due to the promising capabilities of generative models [21, 22, 23, 24]. By coupling the exploratory nature of evolutionary search with the expressive power of generative modeling, GMEAs establish a positive feedback loop between solution generation and model learning. This synergy offers notable potential for solving complex and high-dimensional optimization tasks [4]. However, existing GMEAs are still in their infancy and face three critical challenges: (i) The integration of generative models into evolutionary workflows is often ad hoc and lacks a principled mechanism, thereby necessitating the use of additional heuristic modules. (ii) The training objectives of generative models are frequently borrowed from unrelated domains such as image or text generation, thereby resulting in misalignment with the goals of evolutionary optimization. (iii) Most current methods rely solely on limited population data, which constrains training stability and generalization, particularly in expensive optimization scenarios where evaluation budgets are tight.
These challenges suggest that a fully data-driven evolutionary optimization framework (i.e., one that replaces handcrafted heuristics with end-to-end learning-driven mechanisms) remains an open research goal. In response, we propose Evolutionary Generative Optimization (EvoGO), a data-driven evolutionary optimization framework that eliminates reliance on traditional heuristic operators through generative learning. Our key contributions are summarized as follows:
@@ II-A Non-Data-Driven Approaches
EAs are among the most common non-data-driven approaches for complex optimization. They operate without leveraging data-driven models or learning mechanisms, relying instead on fixed heuristics to explore the search space. These methods can be broadly classified into two categories: (i) those that generate new solutions through handcrafted search operators, and (ii) those that probabilistically model the distribution of promising solutions to guide sampling.
@@
The first category encompasses classic algorithms such as genetic algorithms [9], genetic programming [10], differential evolution [25], and particle swarm optimization [26]. These methods generate new candidate solutions using handcrafted operators like crossover, mutation, or velocity updates. While widely adopted, they rely heavily on domain knowledge and manual tuning, which can limit their flexibility and scalability. To address some of these limitations, the second category i.e, probabilistic modeling-based methods, attempts to guide the search by learning distributions over promising solutions. This category includes two representative families: Evolutionary Strategies (ES) and Estimation of Distribution Algorithms (EDAs).
ES methods emerged from early work on stochastic optimization [11] and have since evolved through several key innovations. Among them, the Covariance Matrix Adaptation ES (CMA-ES) introduced an adaptive mutation mechanism that adjusts the search distribution based on population statistics [27], and was later extended to a fully derandomized formulation [28]. More recently, ES has been applied to large-scale machine learning problems, such as in OpenAI-ES [29], which demonstrated impressive scalability and parallelism in reinforcement learning settings.
In parallel, EDAs integrate statistical modeling into evolutionary search by explicitly learning the distribution of high-quality solutions. The seminal Population-Based Incremental Learning (PBIL) algorithm [30] introduced this idea in discrete spaces, followed by the Univariate Marginal Distribution Algorithm (UMDA) [31], which simplifies modeling by assuming independence among variables. Further developments sought to capture dependencies through more expressive models, such as the extended compact genetic algorithm [32] and the linkage tree genetic algorithm [33].
@@
Despite their conceptual elegance, both ES and EDA methods remain fundamentally non-data-driven in nature. Their update and sampling procedures are still crafted manually and follow rigid algorithmic rules. This restricts their adaptability in diverse or complex scenarios [6, 7]. To enhance flexibility and adaptability, some studies have explored incorporating reinforcement learning (RL) into EAs, aiming to automate the selection or configuration of evolutionary operators [34, 35, 36, 37, 38]. These methods typically employ RL agents to dynamically choose operators or strategies based on past performance. While such integration introduces a level of adaptivity, the learning is still external to the optimization process itself, and often comes with substantial computational overhead and poor sample efficiency [39]. Thus, they do not fundamentally depart from the non-data-driven paradigm, and the search dynamics remain largely heuristic in nature.
@@ II-B Data-Driven Apporoaches
While EAs primarily rely on handcrafted heuristics and static update rules, recent advances have increasingly turned to data-driven approaches that aim to learn from the information accumulated during the optimization process. These approaches leverage machine learning models to improve the efficiency, scalability, or adaptability of evolutionary optimization. Among them, two major categories have emerged: Bayesian Optimization, which constructs surrogate models to guide the search via probabilistic inference, and learnable EAs, which integrate machine learning modules into the evolutionary cycle itself.
@@
Bayesian Optimization (BO) traditionally employs Gaussian processes to model the objective function and constructs acquisition functions to determine promising query points [40]. To improve performance in large-scale and high-dimensional settings, substantial efforts have been made to enhance the scalability of BO [41, 42, 43, 44, 45, 46, 47, 48]. Despite these advances, significant computational overhead persists when BO is applied to complex tasks on modern high-throughput computing platforms.
@@
To address this, researchers have proposed multi-point acquisition strategies to enable batch querying [49, 50, 51, 52]. Among these, Thompson Sampling (TS) [53] has gained popularity for parallel BO [54, 55, 56, 57]. For example, Eriksson et al. integrated TS with a trust region method to achieve scalable BO for both constrained and unconstrained problems, supporting batch sizes of up to 100. Nonetheless, in most cases, the batch sizes remain relatively modest, resulting in underutilization of modern computational resources.
@@
In parallel, Artificial Neural Networks (ANNs) have been explored to improve the batch efficiency of BO [58, 59, 60]. These approaches, however, often repurpose the surrogate model as a binary classifier rather than directly generating new candidate solutions. For instance, Tiao et al. and Oliveira et al. proposed density-ratio-based methods, while Swersky et al. adopted reinforcement learning to train a selection policy for discrete domains. Despite their novelty, these methods still fundamentally rely on direct pointwise optimization for generating new queries [61], which hinders the full scalability of BO.
2) Learnable EAs
Recent advances have witnessed a growing interest in incorporating machine learning techniques into evolutionary algorithms, giving rise to the paradigm of learnable EAs. These methods can be broadly categorized into two main classes: EAs with learnable evaluators, which utilize predictive models to approximate fitness values; and EAs with learnable generators, which aim to improve offspring generation through learned transformations [4].
EAs with Learnable Evaluators
@@
These methods are also known as the Surrogate-Assisted EAs (SAEAs), which are particularly suited for expensive optimization problems where fitness evaluations involve costly simulations. Early frameworks in model-based evolutionary computation [62, 63] established the integration of approximation models to reduce evaluation cost. Traditional SAEAs often employ simple surrogate models such as Radial Basis Functions (RBFs) and Support Vector Machines (SVMs) [64], which are limited in capturing the complexity of high-dimensional landscapes. The introduction of Gaussian Processes (GPs) addressed some of these limitations by enabling uncertainty-aware acquisition functions. For example, Liu et al. [65] combined GPs with EA operators to balance exploration and exploitation. More recent hybrid methods show improved scalability. SAMSO [66] incorporates dual swarm populations with a surrogate model to handle up to 200 decision variables, while AutoSAEA [67] dynamically selects among multiple surrogate models for robust performance across benchmark and real-world tasks. Nonetheless, existing SAEAs often rely heavily on surrogate predictions for candidate selection, which can be computationally intensive and susceptible to model bias [68].
EAs with Learnable Generators
For simplicity, we refer to these methods as Generative Modeling-based Evolutionary Algorithms (GMEAs), which aim to enhance evolutionary optimization by replacing handcrafted variation operators with learned generative mechanisms. These methods have demonstrated promising results across a range of single- and multi-objective problems. Current research in this area predominantly follows two directions: (i) leveraging unsupervised learning for variable grouping and dimensionality reduction [69, 70, 18], and (ii) employing supervised or reinforcement learning to guide the directional improvement of candidate solutions [71, 72, 73, 19, 20]. The first line of work trains generative models to discover latent structure within the search space, often to facilitate more efficient variation. For example, Pan et al. [18] proposed a manifold-learning-based approach that restricts mating operations to low-dimensional subspaces, thereby improving optimization efficiency. The second line focuses on guiding evolutionary search using data-driven feedback. He et al. [73] introduced a GAN-based multi-objective EA where a discriminator distinguishes between real and synthetic populations to train a generator that produces new candidates. Similarly, Liu et al. [19] employed reinforcement learning to adaptively select variation operators based on individual context and historical performance. While GMEAs offer significant potential, they also present several open challenges. These include the difficulty of seamlessly integrating generative models into the evolutionary loop, ambiguities in defining effective training objectives, and a dependence on large population sizes to ensure model generalization and stability.
@@ II-C Discussion
Recent advances in learnable EAs have demonstrated the promise of integrating machine learning into population-based optimization. However, most existing approaches adopt a modular design, treating key components such as fitness prediction, offspring generation, and selection as independent units. While this separation simplifies implementation, it inevitably leads to fragmented learning dynamics and underutilization of the information shared across modules.
In principle, these components are inherently interdependent and can benefit from tighter integration. For example, a shared latent space could act as a unified representation that bridges evaluators and generators: the surrogate model refines this space by encoding fitness-related feedback, while the generator explores it to produce high-potential solutions. Such a design encourages the emergence of a self-reinforcing feedback loop, enabling evaluation and generation to co-adapt based on shared experience, rather than evolving in isolation. Despite this potential, current methods still rely heavily on manually crafted heuristics, rigid coordination rules, and loss functions inherited from unrelated domains (e.g., image synthesis or language modeling). This not only introduces mismatches between local learning objectives and the global goal of optimization, but also undermines the transferability and robustness of the resulting models across different problem settings.
To overcome existing limitations, a promising approach is to develop an end-to-end learning framework that integrates representation learning, solution generation, and fitness evaluation within a cohesive pipeline. Such a framework would ideally (i) eliminate reliance on handcrafted coordination mechanisms, (ii) ensure alignment between the learning objectives of each module and the global optimization goal, and (iii) fully leverage all available information to improve data efficiency and adaptability.

### v2

@@ Abstract
Recent advances in data-driven evolutionary algorithms (EAs) have demonstrated the potential of leveraging historical data to improve optimization accuracy and adaptability. Despite these advancements, existing methods remain reliant on handcrafted process-level operators. In contrast, Evolutionary Generative Optimization (EvoGO) is a fully data-driven framework designed from the objective level, enabling autonomous learning of the entire search process. EvoGO streamlines the evolutionary optimization process into three stages: data preparation, model training, and population generation. The data preparation stage constructs a pairwise dataset to enrich training diversity without incurring additional evaluation costs. During model training, a tailored generative model learns to transform inferior solutions into superior ones. In the population generation stage, EvoGO replaces traditional reproduction operators with a scalable and parallelizable generative mechanism. Extensive experiments on numerical benchmarks, classical control problems, and high-dimensional robotic tasks demonstrate that EvoGO consistently converges within merely 10 generations and substantially outperforms a wide spectrum of optimization approaches, including traditional EAs, Bayesian optimization, and reinforcement learning based methods. Code is available at: https://github.com/EMI-Group/evogo
@@ I Introduction
Many real-world optimization tasks involve objectives that can be accessed only through simulations, legacy software, or physical experiments [1, 2]. Such problems are typically non-analytical, noisy, and gradient-free, making them challenging for conventional optimization techniques. In these scenarios, Evolutionary Algorithms (EAs) provide a robust and widely applicable framework due to their global search capability, insensitivity to differentiability, and ease of parallelization. They have achieved success in diverse fields such as hyperparameter tuning, biomedical engineering, and aerospace design [3]. Nevertheless, while EAs offer a practical starting point for complex black-box problems, their classical forms remain limited in capturing structure or reusing information across generations [4, 5, 6, 7].
To enhance efficiency and stability, subsequent research reformulated EAs from an operator-based to a distribution-based paradigm. Representative examples include Evolution Strategies (ES) [8] and Estimation of Distribution Algorithms (EDAs) [9], which replace handcrafted variation operators with statistical modeling and parameterized sampling. This allows population updates through distributional shifts rather than discrete manipulations, while maintaining desirable invariance properties such as rotation or scale invariance [10]. By estimating gradients in the distribution space, these methods introduce mechanisms like natural gradient ascent, covariance matrix adaptation [11], and factorized or low-rank modeling to better capture dependencies among decision variables[12]. Such strategies effectively adjust the sampling distribution using the data collected from previous evaluations, thereby improving sample efficiency and convergence stability [13]. However, their optimization still focuses on fitting the sampling distribution itself, rather than learning a coherent decision process that couples evaluation and generation in a data-driven manner.
To further enhance the consistency, adaptivity, and data efficiency of evolutionary optimization, subsequent studies have shifted attention from merely updating the sampling distribution to learning the evaluation process itself [14]. This line of research gave rise to data-driven evolutionary frameworks, particularly surrogate-assisted ones, where the objective landscape is approximated through learned evaluators or ranking models trained on historical samples [15, 16]. By replacing direct evaluations with predictive or ordinal surrogates, these methods achieve more consistent and controllable optimization dynamics while significantly reducing computational cost [17, 18, 19]. Fundamentally, the evaluation step becomes a learned component that embeds accumulated experience into the search process, thereby narrowing the gap between evolutionary optimization and data-driven inference. However, their reliance on imperfect surrogates still introduces model bias, and repeated exploitation of surrogate predictions can lead to premature convergence and loss of exploration diversity.
To mitigate these limitations, generative modeling-based evolutionary algorithms (GMEAs) [20, 21, 22] have emerged as a promising extension of data-driven evolutionary computation. Instead of learning to approximate the evaluation process, GMEAs focus on learning the generation process itself, representing the search dynamics through parameterized probabilistic models. By coupling the exploratory nature of evolutionary search with the expressive power of generative modeling, GMEAs establish a positive feedback loop between solution generation and model learning  [23, 24, 25, 26]. GMEAs are particularly suited for complex black-box optimization problems that exhibit latent dependencies, implicit constraints, or simulation-based evaluations, where gradients or analytical forms are unavailable. Through data-driven modeling of the search distribution, GMEAs achieve a more balanced trade-off between exploration and exploitation, effectively alleviating the model bias and premature convergence issues often observed in surrogate-assisted approaches.
However, existing GMEAs are still at an early stage and face several fundamental challenges. First, their integration with evolutionary workflows often relies on manually designed operators and procedural rules, thereby embedding task-specific heuristics into the algorithm rather than achieving automatic coordination between the generative model and the evolutionary process. Second, the training objectives of current generative models are typically inherited from domains such as image or text generation, leading to misalignment with optimization goals. Third, most methods depend on limited population data, which hampers training stability and generalization under tight evaluation budgets. These limitations highlight the need for a fully data-driven evolutionary framework that replaces handcrafted heuristics with end-to-end, learning-based mechanisms. To this end, we propose Evolutionary Generative Optimization (EvoGO), a unified framework that achieves evolutionary optimization entirely through generative learning. Our key contributions are summarized as follows:
@@ II-A Non-Data-Driven Approaches
EAs represent the most prevalent non-data-driven paradigm for complex optimization. These approaches rely on fixed heuristics and do not leverage learned, data-driven models to guide search. EAs can be broadly categorized based on their mechanism: (i) methods that generate solutions via handcrafted search operators, and (ii) methods that probabilistically model the distribution of promising solutions to guide sampling.
@@
Methods based on handcrafted operators, such as genetic algorithms [27], genetic programming [28], differential evolution [29], and particle swarm optimization [30], constitute the first category. These algorithms rely on fixed, manual operators (e.g., crossover or mutation) to explore the search space. Despite their widespread adoption, their performance is heavily dependent on domain knowledge and requires substantial manual tuning, which restricts both flexibility and scalability. To address these limitations, the second category attempts to guide the evolutionary process by learning distributions over promising solutions.
The second category comprises Distribution-based EAs, which aim to improve search efficiency by modeling the distribution of successful solutions. This category includes Evolution Strategies (ES) and Estimation of Distribution Algorithms (EDAs). ES, originating from early stochastic optimization studies [31], achieved a major advancement with Covariance Matrix Adaptation ES (CMA-ES), which adaptively adjusts the sampling distribution based on population statistics [11] and was later extended to a fully derandomized formulation [32]. More recently, ES variants such as OpenAI-ES [33] have demonstrated strong scalability and parallelism in large-scale machine learning tasks. In parallel, EDAs explicitly incorporate statistical modeling into evolutionary search by learning the distribution of promising solutions. Early examples include Population-Based Incremental Learning (PBIL) [13] and the Univariate Marginal Distribution Algorithm (UMDA) [12]. Subsequent methods, such as the extended compact genetic algorithm [34] and the linkage tree genetic algorithm [35], capture variable dependencies through more expressive probabilistic models.
@@
Despite their conceptual elegance, both ES and EDA methods remain fundamentally constrained by the non-data-driven paradigm. The underlying update and sampling procedures are still manually crafted and adhere to rigid algorithmic rules, which severely restricts their adaptability across diverse or complex optimization scenarios [8, 9]. To address this lack of flexibility, some research efforts have explored incorporating reinforcement learning (RL) into EAs to automate the selection or configuration of evolutionary operators [36, 37, 38, 39, 40]. These methods typically employ external RL agents to dynamically choose strategies based on past performance. However, while this integration offers a degree of adaptivity, the learning process is still external to the core optimization loop, often incurring substantial computational overhead and poor sample efficiency [41]. Consequently, these approaches do not fundamentally depart from the heuristic nature of traditional EAs, and the search dynamics remain largely heuristic in nature.
@@ II-B Data-Driven Apporoaches
Traditional EAs primarily rely on handcrafted heuristics and static update rules. However, recent research has increasingly shifted toward data-driven paradigms that exploit information accumulated during the optimization process. These methods employ machine learning techniques to enhance the efficiency, scalability, and adaptability of evolutionary optimization. Consequently, these approaches are generally categorized into three representative categories: Bayesian Optimization (BO), Surrogate-Assisted Evolutionary Algorithms (SAEAs), and Generative Modeling-based Evolutionary Algorithms (GMEAs).
@@
Bayesian Optimization (BO) traditionally employs Gaussian processes to model the objective function and constructs acquisition functions to determine promising query points [42]. Substantial efforts have been made to enhance the scalability of BO to address performance challenges in large-scale and high-dimensional settings [43, 44, 45, 46, 47, 48, 49, 50]. Despite these advances, significant computational overhead persists, especially when BO is applied to complex tasks on modern high-throughput computing platforms.
@@
To address this persistent computational overhead, researchers have proposed multi-point acquisition strategies designed to enable efficient batch querying [51, 52, 53, 54]. Among these approaches, Thompson Sampling (TS) [55] has gained particular popularity for use in parallel BO schemes [56, 57, 58, 59]. For instance, Eriksson et al. integrated TS with a trust region method to achieve scalable BO for constrained and unconstrained problems, supporting batch sizes up to 100. Nevertheless, in the majority of cases, these batch sizes remain relatively modest, thereby resulting in the underutilization of modern computational resources.
@@
In parallel with multi-point methods, Artificial Neural Networks (ANNs) have also been explored to improve the batch efficiency of BO [60, 61, 62]. These approaches, however, often repurpose the surrogate model as a binary classifier rather than directly generating new candidate solutions. For instance, density-ratio-based methods were proposed by Tiao et al. and Oliveira et al., while Swersky et al. adopted reinforcement learning to train a selection policy for discrete domains. Despite their contributions, these methods still fundamentally rely on direct pointwise optimization for generating new queries [63], a mechanism that ultimately hinders the full scalability of BO.
2) Surrogate-Assisted Evolutionary Algorithms
@@
Surrogate-Assisted Evolutionary Algorithms (SAEAs) aim to enhance sample efficiency in optimization problems where fitness evaluations involve costly simulations [64, 7]. Early frameworks in model-based evolutionary computation introduced approximation models to reduce evaluation cost [65, 66]. Traditional SAEAs typically employ simple surrogate models, such as Radial Basis Functions (RBFs) and Support Vector Machines (SVMs) [67]; however, these models are limited in capturing the complexity of high-dimensional and nonlinear search landscapes. The subsequent introduction of Gaussian Processes (GPs) alleviated some of these constraints by incorporating uncertainty estimation into the infill process. For instance, Liu et al. [68] combined GPs with standard EA operators to effectively balance exploration and exploitation.
However, since only a limited number of training samples are available during the optimization process, a single surrogate model often struggles to generalize across heterogeneous problem domains or adapt to different search stages. To address this issue, recent studies have proposed multi-surrogate frameworks that better capture diverse landscape features and mitigate model bias. These frameworks can be broadly classified into two categories: hierarchical-model SAEAs and ensemble-model SAEAs. The hierarchical approach combines global and local models to improve scalability [16, 69, 70], whereas ensemble paradigms integrate multiple predictors to enhance accuracy and reduce overfitting [71, 72]. In parallel, adaptive-model SAEAs have been developed to automatically select appropriate surrogate models or infill criteria according to evolving problem characteristics [18, 19, 73]. Despite these advances, most SAEAs still rely heavily on surrogate predictions for candidate selection, which can be computationally demanding and prone to model bias [73].
3) Generative Modeling-based Evolutionary Algorithms
Generative Modeling-based Evolutionary Algorithms (GMEAs) aim to learn and adapt the sampling distribution from data to guide the evolutionary search process. Current research in this area mainly follows two directions: (i) leveraging unsupervised learning for variable grouping and dimensionality reduction [74, 75, 20], and (ii) employing supervised or reinforcement learning to guide the directional improvement of candidate solutions [76, 77, 78, 21, 22].
The first line of work trains generative models to discover latent structures within the search space, thereby facilitating more efficient variation. For example, Pan et al. [20] proposed a manifold-learning-based approach that restricts mating operations to low-dimensional subspaces, improving optimization efficiency.
The second line focuses on guiding evolutionary search using data-driven feedback. He et al. [78] introduced a GAN-based multi-objective EA in which a discriminator distinguishes real from synthetic populations to train a generator that produces new candidates. Similarly, Liu et al. [21] employed reinforcement learning to adaptively select variation operators based on individual context and historical performance. While GMEAs show strong potential, several open challenges remain. These include their continued reliance on manually designed heuristic components, the difficulty of seamlessly integrating generative models into the evolutionary loop, and ambiguities in defining effective training objectives.
@@ II-C Discussion
The evolution from non-data-driven to data-driven paradigms in evolutionary computation reflects a broader shift toward leveraging information accumulated during optimization. Despite significant progress across BO, SAEA, and GMEAs, each paradigm exhibits distinct advantages and limitations that restrict its general applicability.
BO offers strong theoretical foundations and high sample efficiency but suffers from poor scalability due to the cubic complexity of Gaussian process inference and reliance on pointwise query generation. In contrast, SAEA improves efficiency by introducing learned evaluators to approximate the fitness landscape; however, its focus on local modeling often results in bias and premature convergence. Finally, GMEA emphasizes learning the distribution of promising solutions from data, enabling better structural modeling and diversity maintenance. Nevertheless, current GMEA implementations still rely on manually designed heuristic components and lack effective coupling between the generation and evaluation phases.
To address the existing limitations identified in current approaches, the research imperative is to develop a unified, end-to-end learning framework that seamlessly integrates representation learning, solution generation, and fitness evaluation within a cohesive pipeline. Such a framework must satisfy key design requirements, specifically seeking to (i) eliminate reliance on ad hoc or handcrafted coordination mechanisms, (ii) ensure rigorous alignment between the learning objectives of each internal module and the global optimization goal, and (iii) effectively leverage all available information to maximize data efficiency and adaptability.